# On the uniqueness of compiling graphs under the parity transformation

### v2

@@ 1.1 The parity architecture
Besides recently developed compilation strategies for the parity transformation [28, 19], the encoding procedure has also been further extensively studied in the field of quantum optimization as well as quantum computation. For example in [29], a modular parallelization method for QAOA and parity-encoded optimization problems has been introduced. Furthermore, in [30, 31] novel techniques for realizing quantum devices to perform quantum optimization with neutral atom platforms using the parity architecture have been proposed. As demonstrated in [32], the parity transformation also enables a new universal quantum computing approach which leads to advantages for crucial quantum algorithms such as the quantum Fourier transform. However, a pivotal practical question, directly related to transformation itself and yet to be thoroughly examined, is under which circumstances non-equivalent optimization problems (those that cannot be transformed from one to the other by re-labeling) can or cannot be solved with the same compiled physical layout. Or in other words, how can the set of optimization problems that can be compiled to a given hypergraph under the parity encoding be described and when is the set uniquely determined? Moreover, does there exist a class of physical layouts which do have this uniqueness property? In this paper, we study such questions in more detail and investigate the case where the products in the optimization problems consist of two-body interactions.
@@ 2 Notation and auxiliary results
    $H\sim_{\mathcal{H}}H^{\prime}:\Longleftrightarrow H\text{ is isomorphic to }H^{\prime}$
@@ 2.1 Edge space and cycle space
[…]
@@
    By $\mathcal{B}_{\mathcal{L}_{H}}\coloneqq\left\{B\subset\mathcal{L}_{H}\setminus\left\{\emptyset\right\}\mid\text{$B$ is basis of $\mathcal{L}_{H}$}\right\}$ we denote the set of all bases of $\mathcal{L}_{H}$.
@@ 2.2 The constraint space of hypergraphs
Let $H=(V,E)$ with $V=\left\{1,2,3,4,5\right\}$ and $E=\left\{\left\{1,2\right\},\left\{2,5\right\},\left\{1,3\right\},\left\{1,2,4\right\},\left\{3,4,5\right\}\right\}$. Define $C_{1}\coloneqq\left\{\left\{2,5\right\},\left\{1,3\right\},\left\{1,2,4\right\},\left\{3,4,5\right\}\right\}$ and $C_{2}\coloneqq\left\{\left\{1,2\right\},\left\{2,5\right\},\left\{1,3\right\},\left\{3,4,5\right\}\right\}$. Then we see that $V_{C_{1}}=V$ as well as |$\left|\left\{e\in C_{1}\mid v\in e\right\}\right|=2$ for all $v\in V_{C_{1}}$, which implies that $C_{1}$ is a constraint. However, |$\left|\left\{e\in C_{2}\mid 4\in e\right\}\right|=1$ which shows that $C_{2}$ is not a constraint in $H$. Figure 3 illustrates the two elements $C_{1}$ and $C_{2}$ in the edge space of $H$.
@@
    |$|\{e\in dC_{1}\triangle C_{2}\mid\,$ $v\in e\}|$
    =|$=\left|\left\{e\in dC_{1}\setminus C_{2}\mid v\in e\right\}\right|+\left|\left\{e\in C_{2}\setminus dC_{1}\mid v\in e\right\}\right|$
    =|$=\left|\left\{e\in dC_{1}\mid v\in e\right\}\setminus\left\{e\in C_{2}\mid v\in e\right\}\right|+\left|\left\{e\in C_{2}\mid v\in e\right\}\setminus\left\{e\in dC_{1}\mid v\in e\right\}\right|$
    =|$=\left|\left\{e\in dC_{1}\mid v\in e\right\}\right|+\left|\left\{e\in C_{2}\mid v\in e\right\}\right|-2\left|\left\{e\in dC_{1}\cap C_{2}\mid v\in e\right\}\right|.$
    By assumption, we have |$\left|\left\{e\in dC_{1}\mid v\in e\right\}\right|$ and |$\left|\left\{e\in C_{2}\mid v\in e\right\}\right|$ is even. Thus, |$|\{e\in dC_{1}\triangle C_{2}\mid v\in e\}|$ is even and hence, $dC_{1}\triangle C_{2}\in\mathcal{C}_{H}$. ∎
@@
    |$\left|\left\{e^{\prime}\in C^{\prime}\mid v^{\prime}\in e^{\prime}\right\}\right|\\ =\bigg{|}\biguplus_{\begin{subarray}{c}e^{\prime}\in C^{\prime}\\ v^{\prime}\in e^{\prime}\end{subarray}}\left\{e^{\prime}\right\}\bigg{|}=\bigg{|}\biguplus_{\begin{subarray}{c}e\in C\\ v^{\prime}\in f(e)\end{subarray}}\left\{f(e)\right\}\bigg{|}=\sum_{\begin{subarray}{c}e\in C\\ v^{\prime}\in f(e)\end{subarray}}\left|\left\{f(e)\right\}\right|=\sum_{\begin{subarray}{c}e\in C\\ f^{-1}(v^{\prime})\in e\end{subarray}}\left|\left\{e\right\}\right|\\ =\left|\left\{e\in C\mid f^{-1}(v^{\prime})\in e\right\}\right|$
    and from the definition of $V_{C^{\prime}}$ we see that $f^{-1}(v^{\prime})\in V_{C}$. Therefore, |$\left|\left\{e^{\prime}\in C^{\prime}\mid v^{\prime}\in e^{\prime}\right\}\right|$ is even, which shows that $f_{\mathcal{C}}(\mathcal{C}_{H})\subset\mathcal{C}_{H^{\prime}}$. Next, we define $C\coloneqq\bigcup_{e^{\prime}\in C^{\prime}}\left\{f^{-1}(e^{\prime})\right\}$ for some $C^{\prime}\in\mathcal{C}_{H^{\prime}}$. Using the same calculations as before, we can deduce that $C\in\mathcal{C}_{H}$. Similarly as in the proof of Remark 2.8 (iii), we then obtain $f_{\mathcal{C}}(C)=C^{\prime}$, which shows that $f_{\mathcal{C}}$ is surjective. The linearity and injectivity of $f_{\mathcal{C}}$ follows from the same arguments as in Remark 2.8 (iii).∎
@@
    Similarly to the cycle space, we define $\mathcal{B}_{\mathcal{C}_{H}}$ as the set of all bases of the constraint space of $H$.
As we have seen in Remark 2.12 (ii), the constraint space can be seen as a variant of an extension of the cycle space for hypergraphs. In [35], the definition of cycles for arbitrary hypergraphs has been extended, and it is almost identical to the one as for graphs in Definition 2.4. For hypergraphs, it is additionally assumed that edges in the cycle are pairwise distinct, which is obviously true for edges in a graph. However, using this definition of cycles for hypergraphs and defining the corresponding cycle space as in Definition 2.6 can lead to a space which is not equal to the constraint space as the following example shows.
@@ 2.3 The parity graph mapping
$B_{1}$ =$=\left\{\left\{\left\{1,2\right\},\left\{2,3\right\},\left\{3,4\right\},\left\{4,1\right\}\right\},\left\{\left\{1,2\right\},\left\{2,4\right\},\left\{4,1\right\}\right\}\right\},$
$B_{2}$ =$=\left\{\left\{\left\{1,2\right\},\left\{2,3\right\},\left\{3,4\right\},\left\{4,1\right\}\right\},\left\{\left\{2,3\right\},\left\{3,4\right\},\left\{4,2\right\}\right\}\right\},$
$B_{3}$ =$=\left\{\left\{\left\{1,2\right\},\left\{2,4\right\},\left\{4,1\right\}\right\},\left\{\left\{2,3\right\},\left\{3,4\right\},\left\{4,2\right\}\right\}\right\}\quad\text{and}$
$\mathcal{B}_{\mathcal{C}_{H}}$ =$=\left\{B_{1},B_{2},B_{3}\right\}.$
@@
Let /=$\mathscr{H}\coloneqq\mathcal{H}/{\sim}=\left\{[H]\mid H\in\mathcal{H}\right\}$ and $\mathscr{P}\coloneqq\left\{\mathcal{P}_{H}\mid H\in\mathcal{H}\right\}$ be the set containing every set of compiled hypergraphs. The parity graph mapping is defined as
@@ 3 Uniqueness of the parity graph mapping
$\mathrm{par}|_{\mathscr{D}}^{-1}(\left\{y\right\})\\ =\left\{[H^{\prime}]\in\mathscr{D}\left|\,\exists\,\text{$\ell$ loop labeling of $\mathcal{P}_{B}$}\colon\dim(\mathcal{L}_{G_{\mathcal{P}_{B},\ell}})=\left|B\right|\text{ and $\mathcal{L}_{G_{\mathcal{P}_{B},\ell}}$ is cycle space of $H^{\prime}$}\right.\right\}$ (4)
@@
|$\left|\left\{[H^{\prime}]\in\mathscr{S}\left|\,\exists\,\text{$\ell$ loop labeling of $\mathcal{P}_{B}$}\colon\dim(\mathcal{L}_{G_{\mathcal{P}_{B},\ell}})=\left|B\right|\text{ and $\mathcal{L}_{G_{\mathcal{P}_{B},\ell}}$ is cycle space of $H^{\prime}$}\right.\right\}\right|=1.$
@@
:$\ell\colon\left\{1,\ldots,m\right\}\to\left\{\left\{i,j\right\}\mid i,j\in\mathbb{N},\ i\neq j\right\}\colon v\mapsto\tilde{f}(h(v)),$
@@
$D_{H,B}\coloneqq\\ \left\{[H^{\prime}]\in\mathscr{D}\left|\,\exists\,\text{$\ell$ loop labeling of $\mathcal{P}_{B}$}\colon\dim(\mathcal{L}_{G_{\mathcal{P}_{B},\ell}})=\left|B\right|\text{ and $\mathcal{L}_{G_{\mathcal{P}_{B},\ell}}$ is cycle space of $H^{\prime}$}\right.\right\},$
@@
    In order to prove the other subset relation, we next show that for any two graphs $H_{1},H_{2}$ with $[H_{1}],[H_{2}]\in\mathscr{D}$ and $\mathcal{P}_{H_{1}}=\mathcal{P}_{H_{2}}$, we have $D_{H_{1},B_{1}}=D_{H_{2},B_{2}}$ for all $B_{1}\in\mathcal{B}_{\mathcal{L}_{H_{1}}},B_{2}\in\mathcal{B}_{\mathcal{L}_{H_{2}}}$. Suppose now $B_{1}\in\mathcal{B}_{\mathcal{L}_{H_{1}}}$ and let $B_{2}\in\mathcal{B}_{\mathcal{L}_{H_{2}}}$ such that $[\mathcal{P}_{B_{1}}]=[\mathcal{P}_{B_{2}}]$. Moreover, let $\ell_{1}$ be a loop labeling of $\mathcal{P}_{B_{1}}$ such that $\mathcal{L}_{G_{\mathcal{P}_{B_{1}},\ell_{1}}}$ is the cycle space of some graph $H_{1}^{\prime}=(V_{1}^{\prime},E_{1}^{\prime})$ with $\dim(\mathcal{L}_{G_{\mathcal{P}_{B_{1}},\ell_{1}}})=\left|B_{1}\right|$. Since $\mathcal{P}_{B_{1}}$ is an isomorphic hypergraph to $\mathcal{P}_{B_{2}}$ we can find by Lemma 3.6 a loop labeling $\ell_{2}$ of $\mathcal{P}_{B_{2}}$ such that the induced graph $G_{\mathcal{P}_{B_{1}},\ell_{1}}=(V_{G_{\mathcal{P}_{B_{1}},\ell_{1}}},E_{G_{\mathcal{P}_{B_{1}},\ell_{1}}})$ is equal to the induced graph $G_{\mathcal{P}_{B_{2}},\ell_{2}}=(V_{G_{\mathcal{P}_{B_{2}},\ell_{2}}},E_{G_{\mathcal{P}_{B_{2}},\ell_{2}}})$ with loop labeling $\ell_{2}$. Therefore, we can conclude that
@@
    and thus $[H_{1}^{\prime}]\in D_{H_{2},B_{2}}$. Analogously, we can deduce that $[H_{2}^{\prime}]\in D_{H_{1},B_{1}}$ for an arbitrary element $[H_{2}^{\prime}]\in D_{H_{2},B_{2}}$.
@@
    By definition of $\ell$, we have that $\ell_{k}$ is a loop labeling with $\ell_{k}(e)\in\mathcal{L}_{G_{H_{k},\ell_{k}}}$ for all $e\in E_{k}$. Next, we show the statement for $k=n-1$. Assume by contradiction that $\dim(\mathcal{L}_{G_{H_{n-1},\ell_{n-1}}})\geq n$. Since $e_{n}\setminus\bigcup_{k=1}^{n-1}e_{k}\neq\emptyset$, we have $\dim(\mathcal{L}_{G_{H,\ell}})\geq n+1$, which is a contradiction. Thus, $\dim(\mathcal{L}_{G_{H_{n-1},\ell_{n-1}}})\leq n-1$ and by (i) we have $\dim(\mathcal{L}_{G_{H_{n-1},\ell_{n-1}}})=n-1$. The remaining statement follows from an inductive argument.∎
@@
We show the statement by verifying it for $H_{k}=(V_{k},E_{k})$, where we used the notation in Lemma 3.10 (ii) for $k=1,\ldots,n$. Suppose for $k=1$ we have $\ell_{1}(e_{1})=\ell(e_{1})\not\in L_{G_{H,\ell}}$. Therefore, we can find disjoint sets $\tilde{e}_{1},\tilde{e}_{2}\subset e_{1}$ with $\ell_{1}(\tilde{e}_{l})\in L_{G_{H,\ell}}$ for $l=1,2$. Thus, we observe that $\left\{\ell_{1}(\tilde{e}_{1}),\ell_{1}(\tilde{e}_{2})\right\}$ is linearly independent in $\mathcal{L}_{G_{H_{1},\ell_{1}}}$ and hence $\dim(\mathcal{L}_{G_{H_{1},\ell_{1}}})\geq 2$, which is a contradiction by Lemma 3.10 (ii). Now, assume that statement holds for some $1\leq k<n$ and suppose by contradiction that there exist $\tilde{e}_{1},\tilde{e}_{2}\subset e_{k+1}$ with $\ell(\tilde{e}_{l})\in L_{G_{H,\ell}}$ for $l=1,2$. If $\ell(\tilde{e}_{l})\in\mathcal{L}_{G_{H_{k},\ell_{k}}}$ for some $l\in\left\{1,2\right\}$, then we could find $c_{1},\ldots,c_{k}\in\mathbb{Z}_{2}$ with $\ell(\tilde{e}_{l})=c_{1}\ell(e_{1})\triangle\ldots\triangle c_{k}\ell(e_{k})$. By injectivity of $\ell$, we could derive $\tilde{e}_{l}=c_{1}e_{1}\triangle\ldots\triangle c_{k}e_{k}$, which is a contradiction to the second assumption in (9). This implies that $\ell(\tilde{e}_{l})\not\in\mathcal{L}_{G_{H_{k},\ell_{k}}}$ for $l=1,2$ showing that $\left\{\ell_{k+1}(e_{1}),\ldots,\ell_{k+1}(e_{k}),\ell_{k+1}(\tilde{e}_{1}),\ell_{k+1}(\tilde{e}_{2})\right\}$ is a linearly independent set in $\mathcal{L}_{G_{H_{k+1},\ell_{k+1}}}$ and hence, $\dim(\mathcal{L}_{G_{H_{k+1},\ell_{k+1}}})\geq k+2$, which cannot be true by Lemma 3.10 (ii).∎

### v3

@@ 1.1 The parity architecture
Besides recently developed compilation strategies for the parity transformation [28, 19], the encoding procedure has also been further extensively studied in the field of quantum optimization as well as quantum computation. For example in [29], a modular parallelization method for QAOA and parity-encoded optimization problems has been introduced. Furthermore, in [30, 31] novel techniques for realizing quantum devices to perform quantum optimization with neutral atom platforms using the parity architecture have been proposed. As demonstrated in [32], the parity transformation also enables a new universal quantum computing approach which leads to advantages for crucial quantum algorithms such as the quantum Fourier transform. However, a pivotal practical question, directly related to transformation itself and yet to be thoroughly examined, is under which circumstances non-equivalent optimization problems (those that cannot be transformed into one another by re-labeling) can or cannot be solved with the same compiled physical layout. Or in other words, how can the set of optimization problems that can be compiled to a given hypergraph under the parity encoding be described and when is the set uniquely determined? Moreover, does there exist a class of physical layouts which do have this uniqueness property? In this paper, we study such questions in more detail and investigate the case where the products in the optimization problems consist of two-body interactions.
@@ 2 Notation and auxiliary results
[…]
@@ 2.1 Edge space and cycle space
$\mathcal{L}_{H}=\left\{\left\{\left\{1,2\right\},\left\{2,3\right\},\left\{3,4\right\},\left\{4,1\right\}\right\},\left\{\left\{1,4\right\},\left\{4,3\right\},\left\{3,5\right\},\left\{5,1\right\}\right\},\left\{\left\{1,2\right\},\left\{2,3\right\},\left\{3,5\right\},\left\{5,1\right\}\right\},\emptyset\right\}.$
@@
[…]
@@ 2.2 The constraint space of hypergraphs
Let $H=(V,E)$ with $V=\left\{1,2,3,4,5\right\}$ and $E=\left\{\left\{1,2\right\},\left\{2,5\right\},\left\{1,3\right\},\left\{1,2,4\right\},\left\{3,4,5\right\}\right\}$. Define $C_{1}\coloneqq\left\{\left\{2,5\right\},\left\{1,3\right\},\left\{1,2,4\right\},\left\{3,4,5\right\}\right\}$ and $C_{2}\coloneqq\left\{\left\{1,2\right\},\left\{2,5\right\},\left\{1,3\right\},\left\{3,4,5\right\}\right\}$. Then we see that $V_{C_{1}}=V$ as well as $\left|\left\{e\in C_{1}\mid v\in e\right\}\right|=2$ for all $v\in V_{C_{1}}$, which implies that $C_{1}$ is a constraint. However, $\left|\left\{e\in C_{2}\mid 4\in e\right\}\right|=1$ which shows that $C_{2}$ is not a constraint in $H$. Figure 3 illustrates the two elements $C_{1}$ and $C_{2}$ in the edge space of $H$.
@@
    $|\{e\in dC_{1}\triangle C_{2}\mid\,$ $v\in e\}|$
    $=\left|\left\{e\in dC_{1}\setminus C_{2}\mid v\in e\right\}\right|+\left|\left\{e\in C_{2}\setminus dC_{1}\mid v\in e\right\}\right|$
    $=\left|\left\{e\in dC_{1}\mid v\in e\right\}\setminus\left\{e\in C_{2}\mid v\in e\right\}\right|+\left|\left\{e\in C_{2}\mid v\in e\right\}\setminus\left\{e\in dC_{1}\mid v\in e\right\}\right|$
    $=\left|\left\{e\in dC_{1}\mid v\in e\right\}\right|+\left|\left\{e\in C_{2}\mid v\in e\right\}\right|-2\left|\left\{e\in dC_{1}\cap C_{2}\mid v\in e\right\}\right|.$
    By assumption, we have $\left|\left\{e\in dC_{1}\mid v\in e\right\}\right|$ and $\left|\left\{e\in C_{2}\mid v\in e\right\}\right|$ is even. Thus, $|\{e\in dC_{1}\triangle C_{2}\mid v\in e\}|$ is even and hence, $dC_{1}\triangle C_{2}\in\mathcal{C}_{H}$. ∎
@@
    $\left|\left\{e^{\prime}\in C^{\prime}\mid v^{\prime}\in e^{\prime}\right\}\right|\\ =\bigg|\biguplus_{\begin{subarray}{c}e^{\prime}\in C^{\prime}\\ v^{\prime}\in e^{\prime}\end{subarray}}\left\{e^{\prime}\right\}\bigg|=\bigg|\biguplus_{\begin{subarray}{c}e\in C\\ v^{\prime}\in f(e)\end{subarray}}\left\{f(e)\right\}\bigg|=\sum_{\begin{subarray}{c}e\in C\\ v^{\prime}\in f(e)\end{subarray}}\left|\left\{f(e)\right\}\right|=\sum_{\begin{subarray}{c}e\in C\\ f^{-1}(v^{\prime})\in e\end{subarray}}\left|\left\{e\right\}\right|\\ =\left|\left\{e\in C\mid f^{-1}(v^{\prime})\in e\right\}\right|$
    and from the definition of $V_{C^{\prime}}$ we see that $f^{-1}(v^{\prime})\in V_{C}$. Therefore, $\left|\left\{e^{\prime}\in C^{\prime}\mid v^{\prime}\in e^{\prime}\right\}\right|$ is even, which shows that $f_{\mathcal{C}}(\mathcal{C}_{H})\subset\mathcal{C}_{H^{\prime}}$. Next, we define $C\coloneqq\bigcup_{e^{\prime}\in C^{\prime}}\left\{f^{-1}(e^{\prime})\right\}$ for some $C^{\prime}\in\mathcal{C}_{H^{\prime}}$. Using the same calculations as before, we can deduce that $C\in\mathcal{C}_{H}$. Similarly as in the proof of Remark 2.8 (iii), we then obtain $f_{\mathcal{C}}(C)=C^{\prime}$, which shows that $f_{\mathcal{C}}$ is surjective. The linearity and injectivity of $f_{\mathcal{C}}$ follows from the same arguments as in Remark 2.8 (iii).∎
@@
    Analogously to the cycle space, we define $\mathcal{B}_{\mathcal{C}_{H}}$ as the set of all bases of the constraint space of $H$.
As we can observe in Remark 2.12 (ii), the constraint space can be seen as a variant of an extension of the cycle space for hypergraphs. In [35], the definition of cycles for arbitrary hypergraphs has been extended, and it is almost identical to the one as for graphs in Definition 2.4. For hypergraphs, it is additionally assumed that edges in the cycle are pairwise distinct, which is obviously true for edges in a graph. However, using this definition of cycles for hypergraphs and defining the corresponding cycle space as in Definition 2.6 can lead to a space which is not equal to the constraint space as the following example shows.
@@ 2.3 The parity graph mapping
$B_{1}$ $=\left\{\left\{\left\{1,2\right\},\left\{2,3\right\},\left\{3,4\right\},\left\{4,1\right\}\right\},\left\{\left\{1,2\right\},\left\{2,4\right\},\left\{4,1\right\}\right\}\right\},$
$B_{2}$ $=\left\{\left\{\left\{1,2\right\},\left\{2,3\right\},\left\{3,4\right\},\left\{4,1\right\}\right\},\left\{\left\{2,3\right\},\left\{3,4\right\},\left\{4,2\right\}\right\}\right\},$
$B_{3}$ $=\left\{\left\{\left\{1,2\right\},\left\{2,4\right\},\left\{4,1\right\}\right\},\left\{\left\{2,3\right\},\left\{3,4\right\},\left\{4,2\right\}\right\}\right\}\quad\text{and}$
$\mathcal{B}_{\mathcal{C}_{H}}$ $=\left\{B_{1},B_{2},B_{3}\right\}.$
@@
Let $\mathscr{H}\coloneqq\mathcal{H}/{\sim}=\left\{[H]\mid H\in\mathcal{H}\right\}$ and $\mathscr{P}\coloneqq\left\{\mathcal{P}_{H}\mid H\in\mathcal{H}\right\}$ be the set containing every set of compiled hypergraphs. The parity graph mapping is defined as
@@ 3 Uniqueness of the parity graph mapping
[…]
@@
[…]
@@
$\ell\colon\left\{1,\ldots,m\right\}\to\left\{\left\{i,j\right\}\mid i,j\in\mathbb{N},\ i\neq j\right\}\colon v\mapsto\tilde{f}(h(v)),$
@@
[…]
@@
    In order to prove the other subset relation, we next show that for any two graphs $H_{1},H_{2}$ with $H_{1}],[H_{2}]\in\mathscr{D}$ and $\mathcal{P}_{H_{1}}=\mathcal{P}_{H_{2}}$, we have $D_{H_{1},B_{1}}=D_{H_{2},B_{2}}$ for all $B_{1}\in\mathcal{B}_{\mathcal{L}_{H_{1}}},B_{2}\in\mathcal{B}_{\mathcal{L}_{H_{2}}}$. Suppose now $B_{1}\in\mathcal{B}_{\mathcal{L}_{H_{1}}}$ and let $B_{2}\in\mathcal{B}_{\mathcal{L}_{H_{2}}}$ such that $\mathcal{P}_{B_{1}}]=[\mathcal{P}_{B_{2}}]$. Moreover, let $\ell_{1}$ be a loop labeling of $\mathcal{P}_{B_{1}}$ such that $\mathcal{L}_{G_{\mathcal{P}_{B_{1}},\ell_{1}}}$ is the cycle space of some graph $H_{1}^{\prime}=(V_{1}^{\prime},E_{1}^{\prime})$ with $\dim(\mathcal{L}_{G_{\mathcal{P}_{B_{1}},\ell_{1}}})=\left|B_{1}\right|$. Since $\mathcal{P}_{B_{1}}$ is an isomorphic hypergraph to $\mathcal{P}_{B_{2}}$ we can find by Lemma 3.6 a loop labeling $\ell_{2}$ of $\mathcal{P}_{B_{2}}$ such that the induced graph $G_{\mathcal{P}_{B_{1}},\ell_{1}}=(V_{G_{\mathcal{P}_{B_{1}},\ell_{1}}},E_{G_{\mathcal{P}_{B_{1}},\ell_{1}}})$ is equal to the induced graph $G_{\mathcal{P}_{B_{2}},\ell_{2}}=(V_{G_{\mathcal{P}_{B_{2}},\ell_{2}}},E_{G_{\mathcal{P}_{B_{2}},\ell_{2}}})$ with loop labeling $\ell_{2}$. Therefore, we can conclude that
@@
    and thus $H_{1}^{\prime}]\in D_{H_{2},B_{2}}$. Analogously, we can deduce that $H_{2}^{\prime}]\in D_{H_{1},B_{1}}$ for an arbitrary element $H_{2}^{\prime}]\in D_{H_{2},B_{2}}$.
@@
    By definition of $\ell$, we have that $\ell_{k}$ is a loop labeling with $\ell_{k}(e)\in\mathcal{L}_{G_{H_{k},\ell_{k}}}$ for all $e\in E_{k}$. Next, we show the statement for k=n1k=n-1. Assume by contradiction that $\dim(\mathcal{L}_{G_{H_{n-1},\ell_{n-1}}})\geq n$. Since $e_{n}\setminus\bigcup_{k=1}^{n-1}e_{k}\neq\emptyset$, we have $\dim(\mathcal{L}_{G_{H,\ell}})\geq n+1$, which is a contradiction. Thus, $\dim(\mathcal{L}_{G_{H_{n-1},\ell_{n-1}}})\leq n-1$ and by (i) we have $\dim(\mathcal{L}_{G_{H_{n-1},\ell_{n-1}}})=n-1$. The remaining statement follows from an inductive argument.∎
@@
We show the statement by verifying it for $H_{k}=(V_{k},E_{k})$, where we used the notation in Lemma 3.10 (ii) for $k=1,\ldots,n$. Suppose for $k=1$ we have $\ell_{1}(e_{1})=\ell(e_{1})\not\in L_{G_{H,\ell}}$. Therefore, we can find disjoint sets $\tilde{e}_{1},\tilde{e}_{2}\subset e_{1}$ with $\ell_{1}(\tilde{e}_{l})\in L_{G_{H,\ell}}$ for $l=1,2$. Thus, we observe that $\left\{\ell_{1}(\tilde{e}_{1}),\ell_{1}(\tilde{e}_{2})\right\}$ is linearly independent in $\mathcal{L}_{G_{H_{1},\ell_{1}}}$ and hence $\dim(\mathcal{L}_{G_{H_{1},\ell_{1}}})\geq 2$, which is a contradiction by Lemma 3.10 (ii). Now, assume that statement holds for some $1\leq k<n$ and suppose that there exist $\tilde{e}_{1},\tilde{e}_{2}\subset e_{k+1}$ with $\ell(\tilde{e}_{l})\in L_{G_{H,\ell}}$ for $l=1,2$. If $\ell(\tilde{e}_{l})\in\mathcal{L}_{G_{H_{k},\ell_{k}}}$ for some $l\in\left\{1,2\right\}$, then we could find $c_{1},\ldots,c_{k}\in\mathbb{Z}_{2}$ with $\ell(\tilde{e}_{l})=c_{1}\ell(e_{1})\triangle\ldots\triangle c_{k}\ell(e_{k})$. By injectivity of $\ell$, we could derive $\tilde{e}_{l}=c_{1}e_{1}\triangle\ldots\triangle c_{k}e_{k}$, which is a contradiction to the second assumption in (9). This implies that $\ell(\tilde{e}_{l})\not\in\mathcal{L}_{G_{H_{k},\ell_{k}}}$ for $l=1,2$ showing that $\left\{\ell_{k+1}(e_{1}),\ldots,\ell_{k+1}(e_{k}),\ell_{k+1}(\tilde{e}_{1}),\ell_{k+1}(\tilde{e}_{2})\right\}$ is a linearly independent set in $\mathcal{L}_{G_{H_{k+1},\ell_{k+1}}}$ and hence, $\dim(\mathcal{L}_{G_{H_{k+1},\ell_{k+1}}})\geq k+2$, which cannot be true by Lemma 3.10 (ii).∎